The AVERAGE length of the pancreas of Gambusia is 34.33mm. Variance is 30.144mm and is following normal distribution. If one Gambusia was selected at random,

1) What is the probability that the pancreas length is between 29 and 32mm? Please answer in context of asymptotic theory (property).

2) What is the probability that the pancreas length is smaller than 30mm?Please answer in context of asymptotic theory (property).
Answer #1

solution Here mean (4)=34-33, S.D() = 30.144 x-4 - 29-34-33 = -0.177 D How x=29, 2= x-1 = 30.1uu - - 32-34-33 = -0.0773 for x